The momentum of a particle is p. The kinetic energy of the particle is doubled.
The momentum is now squre root of 2 p.. kindly explain how to get this
1
Expert's answer
2012-12-25T10:00:03-0500
The momentum of a particle is p. The kineticenergy of the particle is doubled.
The momentum is now square root of 2p.. kindly explain how to get this
p=m*v
p - The momentum of a particle
m- mass of a particle
v- velocity
E=m*v^2/2
E - The kinetic energy
E=m*(p/m)^2/2=p^2/(2m)
p=sqrt(2E*m)
p1=sqrt(2E1*m)
p2=sqrt(2E2*m)
p2/p1=sqrt(E2/E1)
If E2/E1=2
p2=p1*sqrt(2)